# Fractional Infinity Laplacian with Obstacle

### v1

@@ 1. Introduction
The analysis of solutions to the infinity Laplacian equations dates back to the early results of Arronson in [4, 5]. Let $\Omega$ be a Lipschitz domain in $\mathbb{R}^{n}$ and $g$ be a Lipschitz function on $\partial\Omega$. Then, the optimal Lipschitz extension $u$ of the boundary datum $g$ minimizing the $L^{\infty}-$norm of the gradient of $u$ on ${\Omega}$ (i.e. $||\nabla u||_{L^{\infty}(\Omega)}$) is a solution in the viscosity sense of the following Dirichlet infinity Laplacian boundary value problem:
@@
Generalization to the Aronsson Functional $||F(x,u,\nabla u)||_{L^{\infty}(\Omega)}$ has been also extensively studied in [7, 8, 11].
@@
where  $\alpha\in(0,1)$ is fixed,  $s=\alpha-\dfrac{n}{p}$,  $g\in C^{0,\alpha}(\partial\Omega)$  and the fractional Sobolev space $W^{s,p}(\Omega)$ is defined as follows:
@@
$[u]_{s,p,\Omega}=\left(\iint_{\Omega\times\Omega}\dfrac{|u(x)-u(y)|^{p}}{|x-y|^{sp+n}}\,\mathrm{d}x\,\mathrm{d}y\right)^{1/p}.$
@@
Moreover, $u_{\infty}$ is an optimal Hölder extension of the boundary datum $g\in C^{0,\alpha}(\partial\Omega)$, in the sense that the Hölder seminorm $[u_{\infty}]_{\alpha,\Omega}$ is always less than or equal $[u]_{\alpha,\Omega}$ for any $\alpha-$Hölder function $u$ such that $u=g$ on $\partial\Omega$, where
$[u]_{\alpha,\Omega}=\sup_{x,y\in\Omega,\,x\neq y}\dfrac{|u(x)-u(y)|}{|x-y|^{\alpha}}.$
@@
But, the uniqueness of the solution to Problem (1.0.4) in the general nonhomogeneous case (i.e. when $f\neq 0$) is still widely an open question. Moreover, the optimal $C^{0,\alpha}-$regularity of the solution remain open for general functions $f$.
@@
Assume $f:\mathbb{R}\mapsto\mathbb{R}$ is continuous and monotone and the boundary datum $g$ is $\beta-$Hölder for some  $0<\beta<\alpha$. Then, the following fractional infinity Laplacian problem:
$\begin{cases}L[u]=f(u)&\qquad in\,\,\,\,\,\Omega\\ u=g&\qquad on\,\,\,\,\,\partial\Omega\end{cases},$
has a solution $u$. Moreover, $u\in C^{0,\beta}(\overline{\Omega})$.
@@
Assume $f$ is nonnegative, continuous and monotone on $[0,\infty)$ and  $g\in C^{0,\beta}(\partial\Omega)$ non negative with  $0<\beta<\alpha$. Then, there exists a nonnegative $\beta-$Hölder solution $u$ to the following obstacle fractional infinity Laplacian problem:
$\begin{cases}L[u]=f(u)&\qquad in\,\,\,\,\,\{u>0\}\\ u=g&\qquad on\,\,\,\,\,\partial\Omega\end{cases}.$
@@ 2. Preliminaries
Fix $\alpha\in(0,1)$. Then, for all $x,\,y\in\mathbb{R}^{n}$, we have the $\alpha-$triangle inequality:
@@
Let $a,\,b>0$. For any $r\geq 0$, we define the function $h(r)=(r+b)^{\alpha}-r^{\alpha}-b^{\alpha}.$ Notice that
@@
Hence, we infer that $h$ is strictly decreasing on $[0,\infty)$ and so, one has the following inequality:
@@
[…]
@@ 3.1. Subsolutions and Supersolutions
Let $\Omega$ be an open bounded domain, $\alpha\in(0,1]$, and $f:\mathbb{R}\mapsto\mathbb{R}$. We say that $u:\overline{\Omega}\mapsto\mathbb{R}$ is a subsolution (resp. supersolution) to the equation $L[u]=f(u)$ and write $L[u]\geq f(u)$ (resp. $L[u]\leq f(u)$) if and only if $u:\overline{\Omega}\mapsto\mathbb{R}$ is upper semi- continuous (resp. lower semi-continuous), and for any test function $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ such that $u\leq\varphi$ (resp. $u\geq\varphi$) with equality at some $x_{0}\in\Omega$ then $-L[\varphi](x_{0})+f(\varphi(x_{0}))\leq 0$ (resp. $-L[\varphi](x_{0})+f(\varphi(x_{0}))\geq 0$). If the last inequality is strict for every such $\varphi$ and $x_{0}$ we say that $u$ is a strict subsolution (resp. supersolution) and write $L[u]>f(u)$ (resp. $L[u]<f(u)$).
@@
Notice that since $L[-u]=-L[u]$ so if $u$ is a supersolution to $L[u]=f(u)$ then $-u$ is a subsolution to $L[v]=-f(-v)$; this follows from the fact that $-u$ is upper semi-continuous and if $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ is such that $-u\leq\varphi$ with equality at $x_{0}$ then $u-(-\varphi)$ attains a minimum at $x_{0}$ and since $u$ is a supersolution to $L[u]=f(u)$, we get
$-L[-\varphi](x_{0})+f(-\varphi(x_{0}))\geq 0.$
Yet, this implies that $-L[\varphi](x_{0})-f(-\varphi(x_{0}))\leq 0.$
@@
$-L[\varphi_{\delta}](x_{1})+f(\varphi_{\delta}(x_{1}))\leq 0.$
But, $\delta>0$ is arbitrary and $\varphi_{\delta}(x_{1})=\varphi(x_{1})$ so letting $\delta\to 0^{+}$, we infer that $-L[\varphi](x_{1})+f(\varphi(x_{1}))\leq 0$, concluding that $u$ is a subsolution on $\Omega\setminus\{x_{0}\}$.
@@
Assume $f$ is continuous. Let $\mathcal{F}$ be a non-empty family of subsolutions to (1.0.8). Define $v(x):=\sup_{u\in\mathcal{F}}u(x)<\infty$ and assume that $v$ is continuous on $\Omega$. Then, $v$ is a subsolution to (1.0.8).
@@
We show that $-L[v]+f(v)\leq 0$ in the viscosity sense. We proceed by contradiction. Assume there exists $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ such that $v(x)-\varphi(x)\leq 0$ on $\overline{\Omega}$ with equality at $x_{0}$ and such that $-L[\varphi](x_{0})+f(\varphi(x_{0}))>0.$ Take $\varphi_{\delta}(x)=\varphi(x)+\delta|x-x_{0}|^{2}$, with $\delta>0$, a peturbation of $\varphi$. We have  $v(x)\leq\varphi(x)<\varphi_{\delta}(x)$ for $x\neq x_{0}$ and $v(x_{0})=\varphi(x_{0})=\varphi_{\delta}(x_{0})$. Hence, $x_{0}$ is the unique maximum to $v-\varphi_{\delta}$. From Lemma (2.7), we have
$-L[\varphi_{\delta}](x_{0})+f(\varphi_{\delta}(x_{0}))=-L[\varphi_{\delta}](x_{0})+f(\varphi(x_{0}))\geq-L[\varphi](x_{0})+f(\varphi(x_{0}))-4\delta\text{{\rm diam}}(\Omega)^{2-\alpha}>0,$
for  $\delta$  small enough. Then, we deduce the existence of a function that we still call it $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ such that $v<\varphi$ with equality only at $x_{0}$ and such that $-L[\varphi](x_{0})+f(\varphi(x_{0}))>0$.
@@
Let $M_{n}=\sup_{x\in\overline{\Omega}}[u_{n}(x)-\varphi(x)]$, which by upper-semicontinuity of $u_{n}$ and compactness of $\overline{\Omega}$ is attained at some $y_{n}\in\overline{\Omega}$.
Claim.
$M_{n}\to 0$ and $y_{n}\to x_{0}$ as $n\to\infty$.
Proof of the Claim.
Suppose there exists a subsequence $y_{n_{k}}$ that converges to $x_{1}\in\overline{\Omega}$. We have
$M_{n_{k}}=u_{n_{k}}(y_{n_{k}})-\varphi(y_{n_{k}})\leq v(y_{n_{k}})-\varphi(y_{n_{k}})\to v(x_{1})-\varphi(x_{1}).$
On the other hand, by (3.1.1), one has
$M_{n_{k}}\geq u_{n_{k}}(x_{0})-\varphi(x_{0})\geq-\dfrac{1}{n_{k}}.$
Then $\liminf_{k\to\infty}M_{n_{k}}\geq 0$, implying that
$v(x_{1})-\varphi(x_{1})\geq 0,$
but $v<\varphi$ with equality only at $x_{0}$ hence $x_{1}=x_{0}$. We conclude that every convergent subsequence of $y_{n}$ converges to $x_{0}$, and hence by compactness $y_{n}\to x_{0}$ as $n\to\infty$. From the argument above, we get also that
$\lim_{n\to\infty}M_{n}=0.$
Now, we complete the proof of the proposition. We define $\varphi_{n}=\varphi+M_{n}$. Notice that
$u_{n}=(u_{n}-\varphi)+\varphi\leq M_{n}+\varphi=\varphi_{n},$
@@
$0\geq-L[\varphi_{n}](y_{n})+f(\varphi_{n}(y_{n}))\geq-L[M_{n}+\varphi](y_{n})+f(M_{n}+\varphi(y_{n}))=-L[\varphi](y_{n})+f(M_{n}+\varphi(y_{n})).$
Hence, thanks to Lemma (2.5), the claim and the continuity of $f$, we conclude that $-L[\varphi](x_{0})+f(\varphi(x_{0}))\leq 0$, a contradiction. ∎
@@
From Remark 3.2, we obtain a similar stability result for supersolutions, that is, if $\mathcal{G}$ is a family of supersolutions to $L[u]=f(u)$ and if $w(x):=\inf_{u\in\mathcal{G}}u(x)$ is continuous, then $w$ is aslo a supersolution.
@@ 3.2. Perron’s Method
Assume $f$ is non-decreasing and continuous, and $g\in C^{0,\beta}(\partial\Omega)$ for some $\beta>0$. Then, there exist a subsolution $u^{-}$ and a supersolution $u^{+}$ to (1.0.8) such that $u^{-}\leq u^{+}$ on  $\overline{\Omega}$  and  $u^{-}=u^{+}=g$  on  $\partial\Omega$.
@@
Assume without loss of generality that $\beta<\alpha$. We define
$u^{-}(x)=\sup\{g(x_{0})-C\,\psi_{\beta,x_{0}}(x),\,x_{0}\in\partial\Omega\},\qquad\mbox{for all}\,\,\,x\in\overline{\Omega},$
where we recall that $\psi_{\beta,x_{0}}(x)=|x-x_{0}|^{\beta}$ and the constant $C>0$ is to be chosen sufficiently large. For every $x_{0}\in\partial\Omega$, thanks to the monotonicity of $f$, we see that
$-L[g(x_{0})-C\,\psi_{\beta,x_{0}}](x)+f(g(x_{0})-C\,\psi_{\beta,x_{0}}(x))\leq C\,L[\psi_{\beta,x_{0}}](x)+f(||g||_{\infty}).$
@@
$L[\psi_{\beta,x_{0}}](x)\leq|x-x_{0}|^{\beta-\alpha}\left(\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}-1\right)\leq\text{{\rm diam}}(\Omega)^{\beta-\alpha}\left(\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}-1\right).$

### v2

@@ 1. Introduction
The analysis of solutions to the infinity Laplacian equations dates back to the early results of Aronsson in [4, 5]. Let $\Omega$ be a Lipschitz domain in $\mathbb{R}^{n}$ and $g$ be a Lipschitz function on $\partial\Omega$. Then, the optimal Lipschitz extension $u$ of the boundary datum $g$ minimizing the $L^{\infty}$-norm of the gradient of $u$ on $\Omega$} (i.e. $\nabla u||_{L^{\infty}(\Omega)}$) is a solution in the viscosity sense of the following Dirichlet infinity Laplacian boundary value problem:
@@
Generalization to the Aronsson Functional $F(x,u,\nabla u)||_{L^{\infty}(\Omega)}$ has been also extensively studied in [7, 8, 12].
@@
where  $\alpha\in(0,1]$ is fixed,  $s=\alpha-\dfrac{n}{p}$,  $g\in C^{0,\alpha}(\partial\Omega)$  and the fractional Sobolev space $W^{s,p}(\Omega)$ is defined as follows:
@@
$u]_{s,p,\Omega}=\left(\iint_{\Omega\times\Omega}\dfrac{|u(x)-u(y)|^{p}}{|x-y|^{sp+n}}\,\mathrm{d}x\,\mathrm{d}y\right)^{1/p}.$
@@
Moreover, $u_{\infty}$ is an optimal Hölder extension of the boundary datum $g\in C^{0,\alpha}(\partial\Omega)$, in the sense that the Hölder seminorm $u_{\infty}]_{\alpha,\Omega}$ is always less than or equal $u]_{\alpha,\Omega}$ for any $\alpha$-Hölder function $u$ such that $u=g$ on $\partial\Omega$, where
$u]_{\alpha,\Omega}=\sup_{x,y\in\Omega,\,x\neq y}\dfrac{|u(x)-u(y)|}{|x-y|^{\alpha}}.$
@@
The uniqueness of the solution to $L[u]=f(x)$ when $f$ is signed, continuous and bounded on $\Omega$ is studied under some growth condition on the solution outside the domain $\Omega$ in [14], except there the operator $L$ is slightly different where the supremum and the infimum are taken over the whole space $\mathbb{R}^{n}$. The uniqueness in the general nonhomogeneous case Problem (1.0.4) is still widely an open question. Moreover, the optimal $C^{0,\alpha}$-regularity of the solution remain open for general functions $f$.
@@
Assume $f:\mathbb{R}\mapsto\mathbb{R}$ is continuous and monotone non-decreasing, and the boundary datum $g$ is continuous on $\partial\Omega$. Then, the following fractional infinity Laplacian problem:
$\begin{cases}L[u]=f(u)&\qquad in\,\,\,\,\,\Omega,\\ u=g&\qquad on\,\,\,\,\,\partial\Omega.\end{cases}$
has a solution $u$, which is locally $\beta$-Hölderian for any $0<\beta<\alpha\leq 1$. Moreover, $u\in C^{0,\beta}(\overline{\Omega})$ as soon as $g\in C^{0,\beta}(\partial\Omega)$. In addition, $u$ is locally $\alpha$-Hölderian if $f$ is nonnegative, for any $\alpha\in(0,1)$.
@@
Assume  $0<\alpha<1$, $f$ is nonnegative, continuous and monotone non-decreasing on $0,\infty)$ and $g\in C^{0,\beta}(\partial\Omega)$ (for some $0<\beta<\alpha$) is nonnegative. Then, there exists a nonnegative $\beta$-Hölder solution $u$ to the following obstacle fractional infinity Laplacian problem:
$\begin{cases}L[u]=f(u)&\qquad in\,\,\,\,\,\{u>0\},\\ u=g&\qquad on\,\,\,\,\,\partial\Omega.\end{cases}$
Moreover, $u$ is locally $\alpha$-Hölder continuous on $\Omega$.
@@ 2. Preliminaries
Fix $\alpha\in(0,1]$. Then, for all $x,\,y\in\mathbb{R}^{n}$, we have the $\alpha$-triangle inequality:
@@
Let a,b>0a,\,b>0. For any $r\geq 0$, we define the function $h(r)=(r+b)^{\alpha}-r^{\alpha}-b^{\alpha}.$ Notice that
@@
Hence, we infer that $h$ is strictly decreasing on $0,\infty)$ and so, one has the following inequality:
@@
$\begin{cases}0&\,\,\,\mbox{if}\,\,\,\,\,0<\alpha<1,\\ \nabla\varphi(x_{0})\cdot{\bf e}&\,\,\,\mbox{if}\,\,\,\,\,\alpha=1.\end{cases}$
@@ 3.1. Subsolutions and Supersolutions
Let $\Omega$ be an open bounded domain, $\alpha\in(0,1]$, and $f:\mathbb{R}\mapsto\mathbb{R}$. We say that $u:\overline{\Omega}\mapsto\mathbb{R}$ is a subsolution (resp. supersolution) to the equation $L[u]=f(u)$ and write $L[u]\geq f(u)$ (resp. $L[u]\leq f(u)$) if and only if $u:\overline{\Omega}\mapsto\mathbb{R}$ is upper semi- continuous (resp. lower semi-continuous), and for any test function $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ such that $u\leq\varphi$ (resp. $u\geq\varphi$) with equality at some $x_{0}\in\Omega$ then $L[\varphi](x_{0})+f(\varphi(x_{0}))\leq 0$ (resp. $L[\varphi](x_{0})+f(\varphi(x_{0}))\geq 0$). If the last inequality is strict for every such $\varphi$ and $x_{0}$ we say that $u$ is a strict subsolution (resp. supersolution) and write $L[u]>f(u)$ (resp. $L[u]<f(u)$).
@@
Notice that since L[u]=L[u]L[-u]=-L[u] so if $u$ is a supersolution to $L[u]=f(u)$ then u-u is a subsolution to L[v]=f(v)L[v]=-f(-v); this follows from the fact that u-u is upper semi-continuous and if $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ is such that $u\leq\varphi$ with equality at $x_{0}$ then $u-(-\varphi)$ attains a minimum at $x_{0}$ and since $u$ is a supersolution to $L[u]=f(u)$, we get
$L[-\varphi](x_{0})+f(-\varphi(x_{0}))\geq 0.$
Yet, this implies that $L[\varphi](x_{0})-f(-\varphi(x_{0}))\leq 0.$
@@
$L[\varphi_{\delta}](x_{1})+f(\varphi_{\delta}(x_{1}))\leq 0.$
But, $\delta>0$ is arbitrary and $\varphi_{\delta}(x_{1})=\varphi(x_{1})$ so letting $\delta\to 0^{+}$, we infer that $L[\varphi](x_{1})+f(\varphi(x_{1}))\leq 0$, concluding that $u$ is a subsolution on $\Omega\setminus\{x_{0}\}$.
@@
Assume $f$ is continuous. Let $\mathcal{F}$ be a non-empty family of subsolutions to (1.0.8). Define $v(x):=\sup_{u\in\mathcal{F}}u(x)<\infty$ and assume that $v$ is upper semi-continuous on $\overline{\Omega}$. Then, $v$ is a subsolution to (1.0.8).
@@
We show that $L[v]+f(v)\leq 0$ in the viscosity sense. We proceed by contradiction. Assume there exists $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ such that $v(x)-\varphi(x)\leq 0$ on $\overline{\Omega}$ with equality at $x_{0}$ and such that $L[\varphi](x_{0})+f(\varphi(x_{0}))>0.$ Take $\varphi_{\delta}(x)=\varphi(x)+\delta|x-x_{0}|^{2}$, with $\delta>0$, a peturbation of $\varphi$. We have  $v(x)\leq\varphi(x)<\varphi_{\delta}(x)$ for $x\neq x_{0}$ and $v(x_{0})=\varphi(x_{0})=\varphi_{\delta}(x_{0})$. Hence, $x_{0}$ is the unique maximum to $v-\varphi_{\delta}$. From Lemma (2.7), we have
$L[\varphi_{\delta}](x_{0})+f(\varphi_{\delta}(x_{0}))=-L[\varphi_{\delta}](x_{0})+f(\varphi(x_{0}))\geq-L[\varphi](x_{0})+f(\varphi(x_{0}))-4\delta\text{{\rm diam}}(\Omega)^{2-\alpha}>0,$
for  $\delta$  small enough.
@@
Let $M_{n}=\sup_{x\in\overline{\Omega}}[u_{n}(x)-\varphi_{\delta}(x)]$, which by upper semi-continuity of $u_{n}$ and compactness of $\overline{\Omega}$ is attained at some $y_{n}\in\overline{\Omega}$. We have
$M_{n}\to 0\qquad\mbox{and}\qquad y_{n}\to x_{0}\qquad\mbox{as}\qquad n\to\infty.$
Indeed, we clearly have that $u_{n}\leq v\leq\varphi_{\delta}$ on $\overline{\Omega}$ and so, $M_{n}\leq 0$. On the other hand, by (3.1.1) and since $v(x_{0})=\varphi_{\delta}(x_{0})$, one has
$M_{n}\geq u_{n}(x_{0})-\varphi_{\delta}(x_{0})\geq-\dfrac{1}{n}.$
Letting $n\to\infty$, we get that $\lim_{n\to\infty}M_{n}=0$. To show that $y_{n}\to x_{0}$, we notice that
$\dfrac{1}{n}\leq M_{n}=u_{n}(y_{n})-\varphi_{\delta}(y_{n})\leq v(y_{n})-\varphi(y_{n})-\delta|y_{n}-x_{0}|^{2}\leq-\delta|y_{n}-x_{0}|^{2}.$
Hence, $|y_{n}-x_{0}|\leq\dfrac{1}{\sqrt{\delta n}}\rightarrow 0$.
Now, we complete the proof of the proposition. We define $\varphi_{n}=\varphi_{\delta}+M_{n}$. Notice that
$u_{n}=(u_{n}-\varphi_{\delta})+\varphi_{\delta}\leq M_{n}+\varphi_{\delta}=\varphi_{n},$
@@
$0\geq-L[\varphi_{n}](y_{n})+f(\varphi_{n}(y_{n}))\geq-L[M_{n}+\varphi_{\delta}](y_{n})+f(M_{n}+\varphi_{\delta}(y_{n}))=-L[\varphi_{\delta}](y_{n})+f(M_{n}+\varphi_{\delta}(y_{n})).$
Hence, thanks to Lemma (2.5), and the continuity of $f$, we conclude that $L[\varphi_{\delta}](x_{0})+f(\varphi_{\delta}(x_{0}))\leq 0$, a contradiction. ∎
@@
From Remark 3.2, we obtain a similar stability result for supersolutions, that is, if $\mathcal{G}$ is a family of supersolutions to $L[u]=f(u)$ and if $w(x):=\inf_{u\in\mathcal{G}}u(x)$ is lower semi-continuous in $\overline{\Omega}$, then $w$ is also a supersolution.
@@ 3.2. Perron’s Method
Assume $f$ is non-decreasing and continuous, and $g$ is continuous on $\partial\Omega$. Then, there exist a subsolution $u^{-}$ and a supersolution $u^{+}$ to (1.0.8) such that $u^{-}\leq u^{+}$ on  $\overline{\Omega}$  and  $u^{-}=u^{+}=g$  on  $\partial\Omega$.
@@
Fix $\beta<\alpha$. For $x_{0}\in\partial\Omega$, $a\in\mathbb{R}$ and $b\geq 0$, we define the function $\phi_{x_{0},a,b}$ on $\overline{\Omega}$ as follows:
$\phi_{x_{0},a,b}(x)=a-b|x-x_{0}|^{\beta}.$
@@
$L[\phi_{x_{0},a,b}](x)=-b\,L[|x-x_{0}|^{\beta}]\geq-b\,|x-x_{0}|^{\beta-\alpha}\left(\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}-1\right)\geq-b\,\text{{\rm diam}}(\Omega)^{\beta-\alpha}\left(\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}-1\right),$ (3.2.1)
for all $x\in\Omega$. We take the set
$S=\bigg\{(x_{0},a,b)\in\partial\Omega\times\mathbb{R}\times[0,\infty):L[\phi_{x_{0},a,b}(x)]\geq f(\phi_{x_{0},a,b}(x))\,\,\,\,\text{in\,\, $\Omega$},\,\,{\phi_{x_{0},a,b}}\leq g\,\,\,\mbox{on}\,\,\,\partial\Omega\bigg\}.$
Notice that $S\neq\emptyset$. Indeed, using the fact that $f$ is non-decreasing and inequality 3.2.1, $(x_{0},a,b)\in S$ as soon as $x_{0}\in\partial\Omega$,  $a\leq\min g$, and
$b\geq\frac{\mbox{diam}(\Omega)^{\alpha-\beta}\,f(a)}{\bigg[1-\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}\bigg]}.$ (3.2.2)
Notice also that for every $(x_{0},a,b)\in S$, we have $a=\phi_{x_{0},a,b}(x_{0})\leq g(x_{0})\leq\max g$ and so, $\phi_{x_{0},a,b}(x)\leq\max g$, for all $x\in\overline{\Omega}.$ We then define
$u^{-}(x)=\sup_{(x_{0},a,b)\in S}\phi_{x_{0},a,b}(x).$
We clearly have $u^{-}(x)<\infty$. Now, we show that $u^{-}=g$ on $\partial\Omega$. By definition of $S$, we have that $\phi_{x_{0},a,b}\leq g$ on $\partial\Omega$ for every $(x_{0},a,b)\in S$ and so, $u^{-}=\sup_{(x_{0},a,b)\in S}\phi_{x_{0},a,b}\leq g$ on $\partial\Omega$.
Given $\varepsilon>0$, by uniform continuity of $g$ there exists $r>0$ such that for all $x,y\in\partial\Omega$ with |xy|<r|x-y|<r, one has
$|g(x)-g(y)|<\epsilon.$
Fix $x_{0}\in\partial\Omega$. Take $a_{\varepsilon}=g(x_{0})-\varepsilon$  and  $b_{\epsilon}\geq\frac{\mbox{diam}(\Omega)^{\alpha-\beta}\,f(a_{\epsilon})}{\bigg[1-\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}\bigg]}$. So, we have $L[\phi_{x_{0},a_{\epsilon},b_{\epsilon}}]\geq f(\phi_{x_{0},a_{\epsilon},b_{\epsilon}})$. Moreover, assume that
$b_{\epsilon}\geq\dfrac{g(x_{0})-\epsilon-\min g}{r^{\beta}}.$
Then, one has
$\phi_{x_{0},a_{\varepsilon},b_{\varepsilon}}(x)=g(x_{0})-\varepsilon-b_{\varepsilon}|x-x_{0}|^{\beta}\leq g(x)\qquad\mbox{for all}\,\,\,x\in\partial\Omega.$
Hence, $(x_{0},a_{\epsilon},b_{\epsilon})\in S.$ In particular, we deduce that
$u^{-}(x_{0})\geq\phi_{x_{0},a_{\varepsilon},b_{\varepsilon}}(x_{0})=g(x_{0})-\varepsilon.$
Letting $\varepsilon\to 0$, we get that $u^{-}(x_{0})\geq g(x_{0})$, concluding that $u^{-}(x_{0})=g(x_{0}),$ for every $x_{0}\in\partial\Omega$.
We next prove that $u^{-}\in C(\overline{\Omega})$. Take $y_{0}\in\overline{\Omega}$ and sequence $y_{n}\in\overline{\Omega}$ that converges to $y_{0}$. Since $u^{-}$ is the supremum of continuous functions, then it is lower semi-continuous and so,
$u^{-}(y_{0})\leq\liminf_{n\to\infty}u^{-}(y_{n}).$
Let $(x_{n},a_{n},b_{n})\in S$ be such that
$u^{-}(y_{n})-\dfrac{1}{n}\leq\phi_{x_{n},a_{n},b_{n}}(y_{n})\leq u^{-}(y_{n}).$ (3.2.3)
Assume $b_{n}$ has no unbounded subsequence that is $\lim_{n\to\infty}b_{n}=\infty$. Since $a_{n}\leq g(x_{n})$, we have
$u^{-}(y_{n})-\dfrac{1}{n}\leq a_{n}-b_{n}|y_{n}-x_{n}|^{\beta}\leq\max g-b_{n}|y_{n}-x_{n}|^{\beta}.$
So,
$|y_{n}-x_{n}|^{\beta}\leq\dfrac{\max g-\inf u^{-}+\dfrac{1}{n}}{b_{n}}.$
Since $u^{-}$ is lower semi-continuous, so the infimum of $u^{-}$ is finite and then, $y_{n}-x_{n}\to 0$, concluding that $x_{n}\to y_{0}$ and so, $y_{0}\in\partial\Omega.$ We then get
$u^{-}(y_{0})=g(y_{0})=\limsup_{n\to\infty}\,g(x_{n})\geq\limsup_{n\to\infty}a_{n}\geq\limsup_{n\to\infty}\phi_{x_{n},a_{n},b_{n}}(y_{n})\geq\limsup_{n\to\infty}\bigg[u^{-}(y_{n})-\dfrac{1}{n}\bigg].$